Question 6 (1 point) assume: a 150 b .75 | 200 Yf 1600 The recessionary gap is: Select an Option
Answer #1

At the equilibrium Y = C+I+G+NX

= 150+0.75Y+200

= 350+0.75Y

(1-0.25)Y = 350

Y = 350/0.25 = 1400

The recessionary gap is 1600-1400 = 200